OpenStudy (anonymous):

A balloon starts rising from rest from the ground with an acceleration of 1.25 m/s2. After 8s, a stone is released from the balloon. The stone will. (a)cover a distance of 40m (b)have a displacement of 50 m (c)each the ground in 4s (d)begin to move down after being released

OpenStudy (anonymous):

My first question is won't there be two accelerations acting on the baloon

Parth (parthkohli):

Oh... yeah, but you will ignore the acceleration due to gravity.

Parth (parthkohli):

1.25 m/s2 here is the overall adjusted acceleration.

Parth (parthkohli):

After 8 seconds, the balloon AND the stone will have a velocity of 10 m/s in the upwards direction. Do you agree?

Parth (parthkohli):

So when the stone is released, it will move in the upward direction for one second before it starts to move down.

OpenStudy (anonymous):

yes by v=u+at

Parth (parthkohli):

The initial height of the balloon:\[h_i = \dfrac{1}{2} \times 1.25 \times 8^2 = 40 m\]

OpenStudy (anonymous):

So b is correct

OpenStudy (anonymous):

why it will move upward

Parth (parthkohli):

Because of inertia. When you are moving upwards and release a stone, then that stone will also initially move upwards.

Parth (parthkohli):

The distance it moves further upwards after being released can also be figured out by using the formula \(H_{max} = u^2/2g\). Here, \(u = 10\) so \(H_{max} = 5\).

OpenStudy (anonymous):

 this path right?

Parth (parthkohli):